# LightMem: Lightweight and Efficient Memory-Augmented Generation

### v3

@@ 1 Introduction
In contrast, human memory efficiently processes information through a hierarchical system: sensory memory pre-filters stimuli, short-term memory actively integrates and reasons over relevant content, and long-term memory selectively consolidates salient information in sleep time.
@@
Building Lightweight Memory. Inspired by the efficiency and structure of human memory, we introduce LightMem, a lightweight memory architecture designed to minimize redundancy while preserving performance. In particular, LightMem emulates human memory through three key components: (1) A pre-compression sensory memory module that filters redundant or low-value tokens from raw input and buffers the distilled content for downstream processing. This initial filtering step reduces noise before information enters the memory pipeline. (2) A topic-aware short-term memory that leverages semantic and topical similarity to dynamically group related utterances into coherent segments. By adaptively determining segment boundaries based on content instead of fixed window sizes, this module produces more concentrated and meaningful memory units. This not only reduces the frequency of memory construction but also enables more precise and efficient retrieval during inference. (3) A sleep-time update mechanism for long-term memory maintenance. New memory entries are initially stored with timestamps to support immediate (“soft”) updates for real-time responsiveness. Later, during designated offline periods (i.e., “sleep”), the system reorganizes, de-duplicates, and abstracts these entries, resolving inconsistencies and strengthening cross-knowledge connections. Crucially, this decouples expensive memory maintenance from real-time inference, enabling reflective, high-fidelity updates without introducing latency. By systematically filtering, organizing, and consolidating relevant information, LightMem substantially reduces computational overhead and API costs while sustaining accurate, coherent reasoning over extended interactions. We detail each component in §3.
@@ 2.1 Conventional Memory Systems for LLMs
2.2 Atkinson–Shiffrin Human Memory Model
Following the Atkinson–Shiffrin human memory model (Atkinson & Shiffrin, 1968), raw environmental information in human brain is first briefly retained in sensory memory, which enables rapid pre-attentive feature extraction and filtering, effectively serving as a form of pre-compression. The processed input can then enter short-term memory (STM), where information and interaction sequences are preserved for tens of seconds to minutes, supporting secondary filtering and more deliberate processing. In contrast, long-term memory (LTM) provides durable storage and undergoes continuous reorganization through updating, abstraction, and forgetting. Importantly, Rasch & Born (2013) highlight that sleep plays a critical role in this reorganization, as oscillatory activity during sleep facilitates the integration and consolidation of memory systems.
@@ 4 Complexity analysis about LightMem
Baselines
$N(L_{\text{sum-in}}+T+L_{\text{sum-out}})$ $M_{1}R_{1}(L_{\text{up-in}}+L_{\text{up-out}}) NN$ $O(N)$
@@
Datasets & Baseline Methods. We use two well-known datasets, LongMemEval (Wu et al., 2025) (specifically the LongMemEval-S split) and LoCoMo (Maharana et al., 2024) to evaluate memory ability. We compare LightMem against several representative baselines of conversational memory modeling. ① Full Text, ② Naive RAG, ③ LangMem (LangChain, 2025), ④ A-MEM (Xu et al., 2025), ⑤ MemoryOS (Kang et al., 2025), ⑥ Mem0 (Chhikara et al., 2025). In addition, all methods use GPT-4o-mini and Qwen3-30B-A3B-Instruct-2507 as the LLM backbones. Details on dataset, baselines, and experimental settings are provided in the Appendix D.
@@ 5.6 Analysis of Sleep-Time Update
Case Study: Memory Update Mechanism Comparison History1: {’Monday, 2 PM’: User is planning a trip to Tokyo.} History2: {’Monday, 4 PM’: User asks about trains to Kyoto.} Hard Update: Overwrites memory -> "User plans Kyoto trip" .  Tokyo context lost LightMem Soft Update: Appends info -> "Tokyo trip + Kyoto inquiry" Full context preserved
@@ Appendix A Background Details
NaiveRAG
$\text{Raw dialog}\to f_{\text{seg}}()$
$\to\{\text{seg}_{i}\}$
$\to f_{\text{index}}()\to\{\text{emb}_{i}\}$ \
Other
Memory
Systems
$\text{Raw dialog}\to f_{\text{seg}}()$
$\to\{\text{seg}_{i}\}$
$\to f_{\text{sum/extract}}()\to\{\text{memory entry}_{i}\}$
$\to f_{\text{index}}()\to\{\text{emb}_{i}\}$
$\to f_{\text{retrieve}}()\to\{\text{related entry}_{i}\}$
$\to f_{\text{update}}()$
$\to\{\text{add, delete, update, merge...}\}$
LightMem
$\text{Raw dialog}\to f_{\text{seg}}()$
$\to\{\text{seg}_{i}\}$
$\to f_{\text{pre\_compress}}()$
$\to\{\text{comp\_seg}_{i}\}$
$\to\text{sensory buffer full}\to$
$f_{\text{topic}}()\to$
$\{\text{topic-wise comp\_seg}_{i}\}$
$\to\text{STM buffer full}$ $\to f_{\text{sum/extract}}()$
$\to\{\text{topic}_{i},\{\text{memory entry}_{j}\}\}$
$\to f_{\text{index}}()\to\{\text{topic}_{i},\{\text{emb}_{j}\}\}$
Offline update trigger
$\{\text{every entry}_{i}\}\to f_{\text{retrieve}}()$
$\to\{\text{related entry}_{j}\}\to\{\text{update queue}\}$
All update queues established
$\to\text{parallel }f_{\text{update}}()$
$\to\{\text{add, delete, update, merge...}\}$
@@ C.2 Category-wise Accuracy
Method
Temporal
( n=133n{=}133)
Multi-Session
(n=133n{=}133)
Knowledge-Update
(n=78n{=}78)
Single-User
(n=70n{=}70)
Single-Assistant
(n=56n{=}56)
Single-Preference
(n=30n{=}30)

### v4

@@ 1 Introduction
Building Lightweight Memory. Inspired by the efficiency and structure of human memory, we introduce LightMem. In particular, LightMem emulates human memory through three key components: (1) A pre-compression sensory memory module that filters redundant or low-value tokens from raw input and buffers the distilled content. (2) A topic-aware short-term memory that leverages semantic and topical similarity to dynamically group related utterances into coherent segments. By adaptively determining segment boundaries based on content instead of fixed window sizes, this module produces more concentrated and meaningful memory units. This not only reduces the frequency of memory construction but also enables more precise and efficient retrieval during inference. (3) A sleep-time update mechanism for long-term memory maintenance. New memory entries are initially stored with timestamps to support immediate (“soft”) updates for real-time responsiveness. Later, during designated offline periods (i.e., “sleep”), the system reorganizes, de-duplicates, and abstracts these entries, resolving inconsistencies and strengthening cross-knowledge connections. Crucially, this decouples expensive memory maintenance from real-time inference, enabling reflective, high-fidelity updates without introducing latency. By systematically filtering, organizing, and consolidating relevant information, LightMem substantially reduces computational overhead and API costs while sustaining accurate, coherent reasoning over extended interactions. We detail each component in §3.
@@ 4 Complexity analysis about LightMem
Baselines $N(L_{\text{sum-in}}+T+L_{\text{sum-out}})$ $M_{1}R_{1}(L_{\text{up-in}}+L_{\text{up-out}}) NN$ $O(N)$
@@
[Uncaptioned image] GLM-4.6
FullText 36.71 103.38
NaiveRAG 73.20 53,725.15
LangMem 49.20 3,052.42 7.03 3,059.45 314.61 5,577.91
A-MEM 70.60 444.95 63.40 1,992.04 403.39 2,903.78 450.40 8,068.80
LightMem
rr=0.5, thth=256 73.00 16.55 14.06 30.61 10.78 1,014.37
rr=0.6, thth=256 73.20 16.55 13.99 30.54 10.78 1,077.69
rr=0.7, thth=512 72.80 16.55 13.91 30.46 10.78 1,038.19
@@
Datasets & Baseline Methods. We use two well-known datasets, LongMemEval (Wu et al., 2025) (specifically the LongMemEval-S split) and LoCoMo (Maharana et al., 2024) to evaluate memory ability. We compare LightMem against several representative baselines of conversational memory modeling. ① Full Text, ② Naive RAG, ③ LangMem (LangChain, 2025), ④ A-MEM (Xu et al., 2025), ⑤ MemoryOS (Kang et al., 2025), ⑥ Mem0 (Chhikara et al., 2025). In addition, all methods use GPT-4o-mini, Qwen3-30B-A3B-Instruct-2507 and GLM-4.6 as the LLM backbones. Details on dataset, baselines, and experimental settings are provided in the Appendix D.
@@ 5.6 Analysis of Sleep-Time Update
Case Study: Memory Update Mechanism Comparison History1: {’Monday, 2 PM’: User is planning a trip to Tokyo.} History2: {’Monday, 4 PM’: User asks about trains to Kyoto.} Hard Update: Overwrites memory -> "User plans Kyoto trip" Tokyo context lost LightMem Soft Update: Appends info -> "Tokyo trip + Kyoto inquiry" Full context preserved
@@
Acknowledgments
We would like to express our sincere gratitude to the anonymous reviewers for their thoughtful and constructive feedback. This work was supported by the National Natural Science Foundation of China (No. 62576307, No. NSFCU23B2055, No. NSFCU19B2027), the Fundamental Research Funds for the Central Universities (226-2023-00138), the 2025 Zhejiang Provincial Center for Disease Control and Prevention Science and Technology Talent Incubation Project (No. 2025-A-04), the 2025 Zhejiang Health Informatics Association Scientific Research Program (Key Project, No. 2025XHZN-Z01), titled “Research on Monitoring and Early Warning Methods of AI Large Model and Infectious Disease Epidemic Data Fusion”, undertaken by the Zhejiang Provincial Center for Disease Control and Prevention, Yongjiang Talent Introduction Programme (2021A-156-G), and Information Technology Center and State Key Lab of CAD&CG, Zhejiang University.
@@ Appendix A Background Details
NaiveRAG $\text{Raw dialog}\to f_{\text{seg}}()$ $\to\{\text{seg}_{i}\}$ $\to f_{\text{index}}()\to\{\text{emb}_{i}\}$ \
Other Memory Systems $\text{Raw dialog}\to f_{\text{seg}}()$ $\to\{\text{seg}_{i}\}$ $\to f_{\text{sum/extract}}()\to\{\text{memory entry}_{i}\}$ $\to f_{\text{index}}()\to\{\text{emb}_{i}\}$ $\to f_{\text{retrieve}}()\to\{\text{related entry}_{i}\}$ $\to f_{\text{update}}()$ $\to\{\text{add, delete, update, merge...}\}$
LightMem $\text{Raw dialog}\to f_{\text{seg}}()$ $\to\{\text{seg}_{i}\}$ $\to f_{\text{pre\_compress}}()$ $\to\{\text{comp\_seg}_{i}\}$ $\to\text{sensory buffer full}\to$ $f_{\text{topic}}()\to$ $\{\text{topic-wise comp\_seg}_{i}\}$ $\to\text{STM buffer full}$ $\to f_{\text{sum/extract}}()$ $\to\{\text{topic}_{i},\{\text{memory entry}_{j}\}\}$ $\to f_{\text{index}}()\to\{\text{topic}_{i},\{\text{emb}_{j}\}\}$ Offline update trigger $\{\text{every entry}_{i}\}\to f_{\text{retrieve}}()$ $\to\{\text{related entry}_{j}\}\to\{\text{update queue}\}$ All update queues established $\to\text{parallel }f_{\text{update}}()$ $\to\{\text{add, delete, update, merge...}\}$
@@ C.2 Category-wise Accuracy
Method Temporal ( n=133n{=}133) Multi-Session (n=133n{=}133) Knowledge-Update (n=78n{=}78) Single-User (n=70n{=}70) Single-Assistant (n=56n{=}56) Single-Preference (n=30n{=}30)